OpenStudy (anonymous):

how do I factor a2-25

OpenStudy (anonymous):

(a+5)*(a-5)

OpenStudy (anonymous):

ok, since it has no a in the middle, u need a + and a - in both solutions to cancel out the x. so it would be: (a+5)(a-5)=0 The 5a's would cancel out each other.

OpenStudy (anonymous):

difference of squares

OpenStudy (anonymous):

this is called the difference of two squares since a^2-b^2 = (a-b)(a+b)

OpenStudy (anonymous):

Some1 thank you and PJW Largo great answers

OpenStudy (anonymous):

glad i could help.